OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (–4, 2) and a radius of 9? I got (x + 4)^2 + (y – 2)^2 = 81 I just wanted to confirm whether it was correct or not. I'm pretty sure the signs change making it (-4,2) and the 81 is the diameter or circumference.

OpenStudy (anonymous):

@amistre64

OpenStudy (gorv):

81 =r^2

OpenStudy (gorv):

(x-x1)^2+(y-y1)^2=r^2

OpenStudy (gorv):

@dricboylan

OpenStudy (amistre64):

the general setup of a circle centered at (a,b) with radius r is: (x-a)^2 + (y-b)^2 = r^2 (x+4)^2+ (y-2)^2 = 9^2 -a = 4 -b = -2 (a,b) = (-4,2) r^2 = 9^2 = 81

OpenStudy (amistre64):

so thats fine

OpenStudy (amistre64):

the "signs change" is a good thing to remember :)

OpenStudy (anonymous):

So I was right?

OpenStudy (amistre64):

you were right